# Reversibility in the Seifert-fibered spaces

### v2

@@ 3. Seifert fibered space
Proof. Since $\pi_{1}(M)/\langle h\rangle$ is a Fuchsian group $\Gamma$, all reversible elements in $\Gamma$ are strongly reversible. By [DG1, Theorem 1.1], it is known that an element in a Fuchsian group is reversible if and only if it is an involution or a product of $2$ involutions. By the representation of the group, the involutions are the elements conjugate to any $c_{i}^{\frac{\mu_{i}}{2}}$ where $\mu_{i}$ is even. Hence, the claim follows. ∎
@@
Proof. If an element in $\pi_{1}(M)$ is reversible then it must be free homotopic to its inverse. This free homotopy forms an embedded Klein bottle in $M$. Let $p:\pi_{1}(M)\to\pi_{1}(G)$ be the projection map from $M$ to the orbit space. If there is a homotopy class of a loop $z$ that is generated by the generating elements of $\pi_{1}(M)$ and has a non-trivial part which cuts the fiber $h$ transitively, then $p(z)$ is a loop in the orbit space. So $h$ is reversible iff the image of $p(z)$ under $\phi$ is non-trivial. Since the image of $\phi$ of $z$ depends on image of $\phi$ of generating set elements, this implies the lemma. Alternatively, since $\phi$ is a homomorphism, it is determined by its actions on the generators of $\pi_{1}(G)$. If $\phi$ is non-trivial, there is some generator $x$ of the form $a_{i}$, $b_{i}$, $x_{i}$ or $d_{i}$ such that $\phi(x)=-1$. Then by the presentation of $\pi_{1}(M)$, $xhx^{-1}=h^{\phi(x)}=h^{-1}$, and so $h$ is reversible.
@@
Proof. On the contrary, let a multiple of $c_{i}$ be a reversible element in $\pi_{1}(M)$ with reverser $\rho$. We can consider that $\rho$ is not a power of $h$, since $h$ is contained in the centralizer of $c_{i}$. Then, the reverser $\rho$ is transverse to $h$, and this implies any power of $\rho$ will always be transverse to $h$. Hence $\rho$ has a non-trivial image in the group $\pi_{1}(M)/\langle h\rangle$, denoted by $[\rho]$. Since, from [DG1], every reversible element in a Fuchsian group is strongly reversible, $[\rho]$ must be of order $2$. This implies, $\rho^{2}=h^{k}$ for some $k\in\mathbb{Z}$, which contradicts our assumption. So, there is no such reverser that reverses any other power of $c_{i}$ except for the powers of the $c^{\mu_{i}}_{i}$. ∎
@@
Then, $h^{\beta_{i}}kh^{\beta_{j}}k^{-1}$ will be identity only if $\beta_{i}=\beta_{j}$ and $\phi(k)=-1.$ It will be same if we choose any other reverser for that particular reversible element since all the reversers are in the form of $x^{l}c^{\frac{\mu_{i}}{2}}.$ In the similar way, we can show that $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{-\frac{\mu_{j}}{2}}k^{-1}$ is reversible if $\phi(k)=1$. If $x$ is not reversible in $\pi_{1}(M)/\langle h\rangle$, a reverser of $x$ in $\pi_{1}(M)$, say $\rho$, must be a power of $h$, since the projection to $\pi_{1}(M)/\langle h\rangle$ must be constant. Then in $\pi_{1}(M)/\langle h\rangle$, $[x]=[x^{-1}]$. This implies that $[x^{2}]=1$. Then, $x^{2}$ is a power of $h$, and $x^{2}$ is also reversible with same reverser $\rho$ that $x$ is. Then,
@@ 4. Application to universal extension of Fuchsian groups
Here, $\pi_{1}(M)$ is the fundamental group of $M$, and $F$ is a Fuchsian group or discrete isometry group of $\rm PGL(2,\mathbb{R})$. So, $\pi_{1}(M)$ is then a universal extension of $F$, which can now be seen as $\pi_{1}(M)/\langle h\rangle$. This gives us a new view-point to consider certain groups as fundamental groups of a Seifert-fibered space. Let us consider $F=\rm PSL(2,\mathbb{Z})$ which is presented by $\langle a,b|a^{2},b^{3}\rangle.$ We have already mentioned the presentation of $\pi_{1}(M)$ for a given $M$ in Equation (1) and Equation (1). We can choose an orientable connected $3$-manifold $M$ in such a way that the $\pi_{1}(M)\cong B_{3}$. $B_{3}$ has a presentation as $\langle\sigma_{1}\sigma_{2}|\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}\rangle$. The normal subgroup of $B_{3}$ is $\mathbb{Z}$ and it is generated by $(\sigma_{1}\sigma_{2})^{3}$. Thus, $B_{3}/\langle\sigma_{1}\sigma_{2}\rangle=\langle\sigma_{1}\sigma_{2}\sigma_{1},\sigma_{1}\sigma_{2}|(\sigma_{1}\sigma_{2})^{3},(\sigma_{1}\sigma_{2}\sigma_{1})^{2}\rangle=\rm PSL(2,\mathbb{Z})\cong\Delta(2,3,\infty)$. Choosing the $c_{i}$ and $\beta_{i}$ suitably, we can consider the Seifert-fibered manifold $M$ determined by $(O,o,0|1,(2,1),(3,1))$. The presentation of the fundamental group of $M$ is then
@@
is a Fuchsian group or a fundamental group of non-orientable orbifold surface. So, it acts as an isometry group in ${\bf H}^{2}$. If $B_{n}/\langle h\rangle$ acts in ${\bf H}^{2}$, then $x^{i}yx^{-i}$ and $y$ are powers of the same element $\psi\in B_{n}/\langle h\rangle$, that is, $x^{i}yx^{-i}=\psi^{u}$ and $y=\psi^{v}$ for some integers $u,v$. But in [DG1], we noticed that $x^{i}$ sends fixed points of $y$ to the fixed points of $x^{i}yx^{-i}$. That means either $x^{i}$ fixes the fixed points of $y$, or interchanges them for every $i$. If it fixes the fixed points then that implies $xy=yx.$ Then $x$ is an element in the center of $B_{n}/\langle h\rangle$, so it must be trivial. This is contradiction in our assumption. So, $x^{i}$ must interchange the fixed points of $y$ for every $i$. Then $x^{2}$ must be an involution (orientation preserving or orientation reversing). This implies $n=4$, and thus $x^{2}yx^{-2}=y^{-1}.$ $y^{-1}$ is $\sigma_{3}$ in $B_{4},$ and then $\sigma_{1}\sigma_{3}$ is an element in the center which contradicts our assumption. So, $B_{n}/\langle h\rangle$ will not be the fundamental group of a Seifert-fibered space for any $n>3$. ∎

### v3

@@ 3. Seifert fibered space
Proof. Since $\pi_{1}(M)/\langle h\rangle$ is a Fuchsian group $\Gamma$, all reversible elements in $\Gamma$ are strongly reversible. By [DG, Theorem 1.1], it is known that an element in a Fuchsian group is reversible if and only if it is an involution or a product of $2$ involutions. By the representation of the group, the involutions are the elements conjugate to any $c_{i}^{\frac{\mu_{i}}{2}}$ where $\mu_{i}$ is even. Hence, the claim follows. ∎
@@
Proof. Since $\phi$ is a homomorphism, it is determined by its actions on the generators of $\pi_{1}(G)$. If $\phi$ is non-trivial, there is some generator $x$ of the form $a_{i}$, $b_{i}$, $x_{i}$ or $d_{i}$ such that $\phi(x)=-1$. Then by the presentation of $\pi_{1}(M)$, $xhx^{-1}=h^{\phi(x)}=h^{-1}$, and so $h$ is reversible. ∎
@@
Proof. On the contrary, let a multiple of $c_{i}$ be a reversible element in $\pi_{1}(M)$ with reverser $\rho$. We can consider that $\rho$ is not a power of $h$, since $h$ is contained in the centralizer of $c_{i}$. Then, the reverser $\rho$ is transverse to $h$, and this implies any power of $\rho$ will always be transverse to $h$. Hence $\rho$ has a non-trivial image in the group $\pi_{1}(M)/\langle h\rangle$, denoted by $[\rho]$. Since, from [DG], every reversible element in a Fuchsian group is strongly reversible, $[\rho]$ must be of order $2$. This implies, $\rho^{2}=h^{k}$ for some $k\in\mathbb{Z}$, which contradicts our assumption. So, there is no such reverser that reverses any other power of $c_{i}$ except for the powers of the $c^{\mu_{i}}_{i}$. ∎
@@
Then, $h^{\beta_{i}}kh^{\beta_{j}}k^{-1}$ will be identity only if $\beta_{i}=\beta_{j}$ and $\phi(k)=-1.$ It will be same if we choose any other reverser for that particular reversible element since all the reversers are in the form of $x^{l}c^{\frac{\mu_{i}}{2}}.$ In the similar way, we can show that $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{-\frac{\mu_{j}}{2}}k^{-1}$ is reversible if $\phi(k)=1$.
If $x$ is not reversible in $\pi_{1}(M)/\langle h\rangle$, a reverser of $x$ in $\pi_{1}(M)$, say $\rho$, must be a power of $h$, since the projection to $\pi_{1}(M)/\langle h\rangle$ must be constant. Then in $\pi_{1}(M)/\langle h\rangle$, $[x]=[x^{-1}]$. This implies that $[x^{2}]=1$. Then, $x^{2}$ is a power of $h$, and $x^{2}$ is also reversible with same reverser $\rho$ that $x$ is. Then,
@@ 4. Application to universal extension of Fuchsian groups
Here, $\pi_{1}(M)$ is the fundamental group of $M$, and $F$ is a Fuchsian group or discrete isometry group of $\rm PGL(2,\mathbb{R})$. So, $\pi_{1}(M)$ is then a universal extension of $F$, which can now be seen as $\pi_{1}(M)/\langle h\rangle$. This gives us a new view-point to consider certain groups as fundamental groups of a Seifert-fibered space. Let us consider $F=\rm PSL(2,\mathbb{Z})$ which is presented by $\langle a,b|a^{2},b^{3}\rangle.$ We have already mentioned the presentation of $\pi_{1}(M)$ for a given $M$ in Equation (1) and Equation (1).
We can choose an orientable connected $3$-manifold $M$ in such a way that the $\pi_{1}(M)\cong B_{3}$. $B_{3}$ has a presentation as $\langle\sigma_{1}\sigma_{2}|\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}\rangle$. The normal subgroup of $B_{3}$ is $\mathbb{Z}$ and it is generated by $(\sigma_{1}\sigma_{2})^{3}$. Thus, $B_{3}/\langle\sigma_{1}\sigma_{2}\rangle=\langle\sigma_{1}\sigma_{2}\sigma_{1},\sigma_{1}\sigma_{2}|(\sigma_{1}\sigma_{2})^{3},(\sigma_{1}\sigma_{2}\sigma_{1})^{2}\rangle=\rm PSL(2,\mathbb{Z})\cong\Delta(2,3,\infty)$. Choosing the $c_{i}$ and $\beta_{i}$ suitably, we can consider the Seifert-fibered manifold $M$ determined by $(O,o,0|1,(2,1),(3,1))$. The presentation of the fundamental group of $M$ is then
@@
is a Fuchsian group or a fundamental group of non-orientable orbifold surface. So, it acts as an isometry group in ${\bf H}^{2}$. If $B_{n}/\langle h\rangle$ acts in ${\bf H}^{2}$, then $x^{i}yx^{-i}$ and $y$ are powers of the same element $\psi\in B_{n}/\langle h\rangle$, that is, $x^{i}yx^{-i}=\psi^{u}$ and $y=\psi^{v}$ for some integers $u,v$. But in [DG], we noticed that $x^{i}$ sends fixed points of $y$ to the fixed points of $x^{i}yx^{-i}$. That means either $x^{i}$ fixes the fixed points of $y$, or interchanges them for every $i$. If it fixes the fixed points then that implies $xy=yx.$ Then $x$ is an element in the center of $B_{n}/\langle h\rangle$, so it must be trivial. This is contradiction in our assumption.
So, $x^{i}$ must interchange the fixed points of $y$ for every $i$. Then $x^{2}$ must be an involution (orientation preserving or orientation reversing). This implies $n=4$, and thus $x^{2}yx^{-2}=y^{-1}.$ $y^{-1}$ is $\sigma_{3}$ in $B_{4},$ and then $\sigma_{1}\sigma_{3}$ is an element in the center which contradicts our assumption. So, $B_{n}/\langle h\rangle$ will not be the fundamental group of a Seifert-fibered space for any $n>3$. ∎
@@
5. Generalised 3-Torsion elements in $\pi_{1}(B_{3})$
Definition 5.1.
An element $g$ in the group $G$ is said to be a generalised $n$-torsion element if it satisfies the following condition:
$k_{1}gk_{1}^{-1}k_{2}gk_{2}^{-1}\dots k_{n}gk_{n}^{-1}=Id$
where, $k_{1},k_{2},\dots,k_{n}\in G.$
Note that, a generalised $2$-torsion element is equivalent to a reversible element by definition. This property is invariant over conjugacy and inversion. In this section, we study the generalised $3$-torsion elements in the braid group with $3$-strains $B_{3}$, and provide a sufficient condition for an element to be a $3$-torsion element. To do that, we first check the sufficient conditions for the generalised $3$-torsion elements in $\rm PSL(2,\mathbb{Z})$.
Lemma 5.2.
If $g\in\rm PSL(2,\mathbb{Z})$ is a generalised $3$-torsion element then, $g$ is a product of $3$-torsion elements.
Proof. Let us consider the presentation of $\rm{PSL(2,\mathbb{Z})}=$ $\langle a,b~{}|~{}a^{2},b^{3}\rangle$. In $\rm PSL(2,\mathbb{Z})$, let $g$ be a generalised $3$-torsion element. If we abelianise $\rm PSL(2,\mathbb{Z})$, then $[g]^{3}=Id$. That means, $g$ is a $3$-ordered element or it is a product of $3$-ordered elements in $\rm PSL(2,\mathbb{Z})$. The commutator $[a,b]$ of $a,b$ is also a product of two 3-ordered elements. Then, the reduced form of $g$ is $\Pi_{i=1}^{m}k_{i}[a,b]k_{i}^{-1}\Pi_{j=1}^{n}b_{i}$ where, $b_{i},k_{i}\in\rm PSL(2,\mathbb{Z})$ and $|b_{j}|=3$. ∎
If $g$ is an elliptic element in $\rm PSL(2,\mathbb{Z})$, then it is known to be of order either $2$ or $3$. Thus, by the previous lemma, a generalised $3$-torsion element which is also an elliptic element must be of order $3$, since the product of three ordered elements cannot be an involution.
Therefore, we can now assume that $g\in\rm PSL(2,\mathbb{Z})$ is a non-elliptic generalised $3$-torsion element i.e., there are $h,k\in\rm PSL(2,\mathbb{Z})$ such that
$ghgh^{-1}kgk^{-1}=Id$
and also, $g$ is conjugate to $ab^{k_{1}}ab^{k_{2}}\dots ab^{k_{n}}$ where, $k_{i}=\pm 1$. Since the property of generalised $3$-torsion element is invariant over conjugacy, it is sufficient to consider
(5.1) $g=ab^{k_{1}}ab^{k_{2}}\dots ab^{k_{n}}$
where, $k_{i}=\pm 1$ and $1\leq k_{i}\leq n$.
Lemma 5.3.
In $\rm PSL(2,\mathbb{Z})$, there are only two parabolic elements which are generalised $3$-torsion elements upto conjugacy, $(ab)^{\pm 2}$.
Proof. There is only one type of parabolic conjugacy classes in $\rm PSL(2,\mathbb{Z)}$ i.e., $(ab)^{n}$, where, $n\in\mathbb{Z}$. The matrix form of $(ab)^{n}$ is $\begin{bmatrix}1&&n\\ 0&&1\end{bmatrix}$. Let $g=(ab)^{n}$ be a generalised $3$-torsion element i.e., there exists $h,k\in\rm PSL(2,\mathbb{Z})$ such that,
$ghgh^{-1}kgk^{-1}=Id$
$\Rightarrow ghgh^{-1}=kg^{-1}k^{-1}.$
Thus, $ghgh^{-1}$ is a parabolic element, since any conjugate of $g$ or $g^{-1}$ needs to be parabolic. Therefore, $|trace(ghgh^{-1})|$ is 2. Suppose $h$ has the matrix form $\begin{bmatrix}a&&b\\ c&&d\end{bmatrix}.$ Let the matrix form of $kg^{-1}k^{-1}$ be denoted by $A$. Then,
$A=\begin{bmatrix}1&&n\\ 0&&1\end{bmatrix}\begin{bmatrix}a&&b\\ c&&d\end{bmatrix}\begin{bmatrix}1&&n\\ 0&&1\end{bmatrix}\begin{bmatrix}a&&b\\ c&&d\end{bmatrix}^{-1}$
$=\begin{bmatrix}1&&n\\ 0&&1\end{bmatrix}\begin{bmatrix}a&&b\\ c&&d\end{bmatrix}\begin{bmatrix}1&&n\\ 0&&1\end{bmatrix}\begin{bmatrix}d&&-b\\ -c&&a\end{bmatrix}$
$=\begin{bmatrix}a+nc&&b+nd\\ c&&d\end{bmatrix}\begin{bmatrix}d-cn&&-b+an\\ -c&&a\end{bmatrix}$
$=\begin{bmatrix}1-n^{2}c^{2}-acn&&acn^{2}+a^{2}n+n\\ -c^{2}n&&1+acn\end{bmatrix}.$
Since, $A$ is parabolic, this implies, $|trace(A)|=2.$ The obvious way to get $|trace(A)|=2$ is if we consider $c=0$, which will give us $h=g^{i}$ for some $i\in\mathbb{Z}$. Then $A$ will be $g^{2}$, which cannot be true, since $g^{-1}$ is not conjugate to $g^{2}$. Now, we consider $c\neq 0.$ That means
$|trace(A)|=2$
$\Rightarrow|2-n^{2}c|=2$
$\Rightarrow 2-n^{2}c=-2$
since the $22-n^{2}c=$ 2 case will be reduced to $c=0.$ So,
$n^{2}c=4.$
That implies, either $n=1,\ c=4~{}\hbox{or}~{}n^{2}=4,\ c=1$. Since $g$ cannot be seen as a product of $3$-torsion elements if we take $n=1$, therefore we conclude that $g$ may be one of $(ab)^{\pm 2}.$
Without loss of generality, let us choose $g$ to be $(ab)^{2}.$ This implies, $g=abab.$ Also note that, $ababb^{-1}abab^{-1}baba=Id.$ Therefore, $ghgh^{-1}kgk^{-1}=Id$ where, $h=b^{-1}$ and $k=b.$ This verifies that $g$ is indeed a generalised $3$-torsion element.∎
In the proof of the previous lemma, we have obtained that a parabolic element is the product of two 3-ordered elements. Therefore, let us now check if this is also true for hyperbolic elements.
Lemma 5.4.
Let $g$ be a hyperbolic element in $\rm PSL(2,\mathbb{Z})$ and $g$ can be seen as a product of two 3-ordered elements i.e., $g=b_{1}b_{2}$, where $|b_{1}|=|b_{2}|=3.$ Then, $g$ is a generalised 3-ordered torsion element.
Proof. $g=b_{1}b_{2}$ this implies,
$b_{1}b_{2}b^{-1}_{2}b_{1}b^{-1}_{2}b_{2}b_{1}=Id$
$\Rightarrow gb^{-1}_{2}gb_{2}b_{2}gb^{-1}_{2}=Id.$
Thate means, $ghgh^{-1}kgk^{-1}=Id$ where, $h=b_{2}^{-1}$ and $k=b_{2}.$ Hence, this proves the lemma. ∎
Theorem 5.5.
Suppose $g$ be an element in $\rm PSL(2,\mathbb{Z})$, it is a generalized $3$-torsion element, if any of the followings hold:
  1. (1)
    $g$ is an elliptic element, then $g$ must be order of $3$
  2. (2)
    $g$ is a parabolic element, then $g$ is conjugate to any of $(ab)^{\pm 2}.$
  3. (3)
    $g$ is a hyperbolic element, then $g$ is conjugate to an element which is product of two $3$-ordered elements.
Proof. $(1)$ is easy to see from the above discussion. $(2)$ follows from Lemma 5.3 and $(3)$ follows from Lemma 5.4. ∎
Notice that the condition of the hyperbolic elements obtained in the previous theorem holds true for any other Fuchsian group. This motivates us to see the generalised $3$-torsion elements in the universal central extension of Fuchsian groups, i.e., in Seifert fibered groups.
Proposition 5.6.
Let us consider a Seifert-fibered space $M$ with fundamental group (1) with some elements $c_{1}$ and $c_{2}$ such that $c_{1}^{3p}=h^{\mu_{1}}$ and $c_{2}^{3q}=h^{\mu_{2}}$ for some $p,q\in\mathbb{N}$. Then $e_{1}^{p}e_{2}^{q}h^{x}$ is a generalised $3$-torsion element if $03x+\mu_{1}+\mu_{2}=$ 0, where $e_{i}$ are conjugate to $c_{i}$.
Proof. Consider two elements $c_{1}$ and $c_{2}$ in $M$ where $c_{1}^{3p}=h^{\mu_{1}}$ and $c_{2}^{3q}=h^{\mu_{2}}$ for some $p,q\in\mathbb{N}$ and $\mu_{1},\mu_{2}\in\mathbb{Z}$. We know that in the Fuchsian group $\pi_{1}({M})/\langle h\rangle$, the product of two $3$-ordered elements are generalised $3$-torsion elements.
Now we take one of the lifts of $e_{1}^{p}e_{2}^{q}$ in $\pi_{1}(M)$, where $e_{i}$ conjugate to $c_{i}$ for $i=1,2$, and let $g=e_{1}^{p}e_{2}^{q}h^{x}$, where $x\in\mathbb{Z}.$ Consider $h_{1}=e_{2}^{-q}$ and $k=e_{2}^{q}$. If $g$ satisfies the condition
$gh_{1}gh_{1}^{-1}kgk^{-1}=Id,$
then
$e_{1}^{p}e_{2}^{q}h^{x}e_{2}^{-q}e_{1}^{p}e_{2}^{q}h^{x}e_{2}^{q}e_{2}^{q}e_{1}^{p}e_{2}^{q}h^{x}e_{2}^{-q}=Id$
$\Rightarrow e_{1}^{p}h^{x}e_{1}^{p}h^{x}e_{2}^{3q}e_{1}^{p}h^{x}=Id$
$\Rightarrow e_{1}^{3p}h^{3x+\mu_{2}}=Id.$
Since, $h$ has infinite order in $\pi(M)$, $03x+\mu_{1}+\mu_{2}=$ 0 . In that case, $g$ is a generalised $3$-torsion element in $\pi_{1}(M)$. ∎
Remark 5.7.
We can find some sufficient conditions for the existence of generalised $n$-torsion elements in a Seifert-fibered space following similar arguments as above. In particular, if there exist elements $c_{i}$ for $i=1,2$ such that $c_{i}^{np_{i}}=h^{\mu_{i}}$ in the fundamental group, then any conjugate of $c_{1}^{p_{1}}c_{2}^{p_{2}}h^{x}$ is a generalised $n$-torsion element if $nx+\mu_{1}+\mu_{2}=0$.
In the particular case of the braid group $B_{3}$, we now have the following corollary.
Corollary 5.8.
Consider $M=S^{3}-K$ where $K$ is the trefoil knot, and hence $\pi_{1}(M)=B_{3}=\langle c_{1},c_{2},h|c_{1}^{2}=c_{2}^{3}=h,c_{i}h=hc_{i}\rangle$. Then the elements which are conjugate to $e_{i}e_{j}^{2}h^{-1}$, where $e_{i}$’s are distinct and conjugate to $c_{2}$, are generalised $3$-torsion elements.
Proof. Let us choose two $3$-ordered elements $[e_{1}],[e_{2}]$ in the quotient group $B_{3}/\langle h\rangle$, conjugate to $[c_{2}].$ So, in $B_{3}$, $e_{1}^{3}=h=e_{2}^{3}.$ In $\rm PSL(2,\mathbb{Z})$ we have noticed that the product of two $3$-ordered elements are generalised $3$-torsion elements.
So, we have two choices of generalised 3-torsion elements. For the first case, we take one of the lifts of $e_{1}e_{2}$, say $g=e_{1}e_{2}h^{x}$, where $x\in\mathbb{Z}$. Let g satisfy the condition
$gh_{1}gh_{1}^{-1}kgk^{-1}=Id,$
and let us choose $h_{1}=e_{2}^{-1}$ and $k=e_{2}$. Then,
$e_{1}e_{2}h^{x}e_{2}^{-1}e_{1}e_{2}h^{x}e_{2}e_{2}e_{1}e_{2}h^{x}e_{2}^{-1}=Id$
$\Rightarrow e_{1}h^{x}e_{1}h^{x}e_{2}^{3}e_{1}h^{x}=Id$
$\Rightarrow h^{3x+2}=Id.$
This case is not possible, since $h$ has infinite order and there is no integer $x$ such that $03x+2=$ 0 . For the next case we may consider $g$ as a lift of $e_{1}e_{2}^{2}$, i.e., $g=e_{1}e_{2}^{2}h^{x}.$ Now, if we continue the same process as previous case, we obtain $h^{3x+3}=Id.$ Therefore, for $x=-1$, $g$ will be a generalised $3$-torsion element. ∎